### v3

# Unveiling Group-Specific Distributed Concept Drift: A Fairness Imperative in Federated Learning

Abstract
In the evolving field of machine learning, ensuring fairness has become a critical concern, prompting the development of algorithms designed to mitigate bias in decision-making processes. However, achieving fairness in the presence of group-specific concept drift remains an unexplored frontier, and our research represents pioneering efforts in this regard. Group-specific concept drift refers to situations where one group experiences concept drift over time while another does not, leading to a decrease in fairness even if accuracy remains fairly stable. Within the framework of Federated Learning, where clients collaboratively train models, its distributed nature further amplifies these challenges since each client can experience group-specific concept drift independently while still sharing the same underlying concept, creating a complex and dynamic environment for maintaining fairness. The most significant contribution of our research is the formalization and introduction of the problem of group-specific concept drift and its distributed counterpart, shedding light on its critical importance in the realm of fairness. Additionally, leveraging insights from prior research, we adapt an existing distributed concept drift adaptation algorithm to tackle group-specific distributed concept drift which uses a multi-model approach, a local group-specific drift detection mechanism, and continuous clustering of models over time. The findings from our experiments highlight the importance of addressing group-specific concept drift and its distributed counterpart to advance fairness in machine learning.
Index Terms:
Fairness, Concept Drift, Federated Learning.
I Introduction
Fairness is a fundamental concern in machine learning, aiming to prevent discrimination in algorithmic decision-making processes with respect to sensitive attributes, such as race or gender [1, 2]. However, achieving fairness becomes even more challenging in the presence of concept drift, which refers to the phenomenon where the underlying data distribution changes over time. In this paper, we first introduce and formalize the concept of group-specific concept drift, which focuses on scenarios where one group, characterized by a sensitive attribute, experiences changes in the data distribution while another group remains unaffected.
To illustrate the importance of the problem, consider an example of group-specific concept drift in the context of loan lending. Suppose a financial institution operates in a region where historically marginalized communities have limited access to financial services. Over time, societal changes, government initiatives, and increased awareness promote financial inclusion, leading to an increase in loan applications from these previously underserved communities. In this evolving environment, group-specific concept drift occurs as the data distribution for loan applicants from marginalized communities changes. Traditional lending models may not adequately capture the creditworthiness and financial profiles of these communities due to the historical underrepresentation and lack of relevant data. Consequently, there is a risk of perpetuating bias in loan approval decisions. Addressing group-specific concept drift is crucial to ensure fair access to loans for individuals from historically marginalized communities.
Federated Learning (FL), a decentralized approach to machine learning, provides a unique setting for studying fairness under group-specific distributed concept drift. In this framework, multiple clients collaboratively train models using their local data, with a central server aggregating their updates to create a global model, preserving client data privacy and sharing only model updates. [3]. However, the distributed nature of FL, combined with the presence of group-specific concept drift, poses new challenges that have not been addressed in the context of fairness. In this context, group-specific concept drift could manifest in specific clients at different times, creating challenges in maintaining fairness and accuracy. As such, in this work, we also introduce and formalize the concept of group-specific distributed concept drift in distributed environments.
To illustrate the importance of the problem of group-specific distributed concept drift, consider several examples across different domains:
Example 1: Going back to the previous example of loan lending, now in a FL setting, consider a scenario where clients representing different regions or demographic groups experience varying rates of change in the data distribution of loan applicants. For instance, a financial institution operating in a diverse geographic landscape may have multiple branches, each serving a unique community with its distinct characteristics. In this FL framework, if one branch undergoes changes in the data distribution for loan applicants from marginalized communities, while another branch remains unaffected, the result is a manifestation of group-specific distributed concept drift. The decentralized nature of FL implies that these shifts in data distribution occur independently across different clients at various times, introducing a layer of complexity in maintaining fairness and accuracy. As the financial landscape evolves and societal changes unfold, addressing group-specific distributed concept drift becomes imperative to uphold fairness and prevent perpetuation of biases in loan approval decisions for diverse and historically underserved communities.
Example 2: Consider an international online fashion retailer using FL to tailor product recommendations for users across different countries. The retailer’s user base encompasses a diverse array of demographic groups, each with distinct fashion inclinations shaped by cultural norms, climate, and socioeconomic factors. Envision a scenario where a particular fashion trend emerges in Country A (i.e. client A) amongst women (i.e. sensitive group - gender) due to cultural phenomena or seasonal shifts. However, this same trend might not resonate or even be visible in Country B (i.e. client B) due to differing cultural norms or consumer preferences. In some instances, what is considered fashionable in Country A might only gain traction among women in Country B after a significant delay, or it might not find relevance at all due to cultural disparities. Consequently, the online fashion retailer encounters group-specific distributed concept drift, wherein the data distribution of user preferences and purchasing behaviors diverges across different countries and within specific demographic segments. While women in Country A may eagerly adopt the latest trend, other regions such as Country B might exhibit different fashion preferences or remain indifferent to the trend altogether. This concept drift presents notable challenges for the retailer’s FL system. Failing to address this group-specific distributed concept drift could result in suboptimal recommendations, diminished user engagement, and the risk of biased or discriminatory outcomes, particularly if certain demographic segments are underrepresented in the recommendation process.
Example 3: Consider a healthcare system serving patients from diverse ages, using FL to develop predictive models for disease diagnosis and treatment recommendations. However, the healthcare landscape is dynamic, with demographic groups experiencing fluctuations in health behaviors, healthcare access, and disease prevalence. Consider a scenario where certain hospitals, operating as individual clients within the FL framework, introduce new drugs or treatment protocols. These interventions aim to address specific health conditions prevalent within their patient populations. However, the introduction of these drugs may inadvertently lead side effects within these hospitals’ patient populations. For instance, the administration of certain drugs for managing a disease may lead to side effects among older patients (i.e. sensitive group - age). Consequently, the data distribution of disease patterns within these hospitals undergoes a noticeable shift, primarily affecting the older demographic. In contrast, younger patients within these hospitals who are less susceptible to developing these side effects as a result of the drug, may not experience the same level of change in disease manifestations. This distributed concept drift primarily affects the older demographic, underscoring the importance of adapting predictive models to evolving data distributions. Addressing this group-specific distributed concept drift is essential to ensure accurate and equitable healthcare outcomes across diverse patient populations, thereby minimizing disparities in healthcare delivery.
Contributions
Our contributions span three key dimensions, each enhancing our understanding and addressing challenges related to group-specific concept drift in FL:
(1) Formalization of group-specific concept drift and its distributed counterpart
We pioneer the introduction and formalization of the concept of group-specific concept drift and group-specific distributed concept drift, shedding light on previously unexplored aspects within the fairness landscape
(2) Experimental framework for studying group-specific distributed concept drift
We establish an experimental framework to study this new challenge, leveraging diverse datasets with varying degrees of sensitive group imbalance. This robust setup provides a foundation for exploring the intricacies of group-specific distributed concept drift.
(3) Algorithm for handling group-specific distributed concept drift
Building upon techniques from prior studies, we propose FairFedDrift: a continuous multi-model clustering approach with a local group-specific drift detection mechanism to effectively address the intricate problem of group-specific distributed concept drift. Our experimental results demonstrate the efficacy of FairFedDrift in effectively detecting and managing group-specific distributed concept drift in FL.
The remainder of this work is structured as follows: Section 2 provides some background and related work on the subject, Section 3 describes the problem statement, Section 4 presents the proposed approach, Section 5 describes the methodology and experimental design, Section 6 presents the results and discussion, Section 7 presents the conclusions and possible future research.
II Related Work
II-A Fairness
Fairness-aware machine learning has gained significant attention, and algorithms designed to promote fairness can be categorized into three groups based on the stage at which they are implemented: pre-processing, in-processing, and post-processing [1, 2, 4, 5]. The algorithm proposed in this study belongs to the in-processing category. Fairness metrics can be broadly classified into two main groups: group fairness and individual fairness. Group fairness emphasizes that the outcomes of an algorithm do not disproportionately favor or harm any particular group based on characteristics such as race, gender, ethnicity, religion, or socioeconomic status. On the other hand, individual fairness aims to ensure similar predictions for individuals who share similar characteristics [1, 4]. This research specifically focuses on optimizing group fairness.
II-B Fair Federated Learning
In the realm of FL, there has been a growing focus on achieving fairness [6, 7, 8, 9, 10, 11]. These methods aim to mitigate bias while preserving privacy in a decentralized environment using techniques such as fair global model aggregation techniques [6, 7, 8, 9] and local debiasing methodologies at each client [10, 7].
These approaches, designed for static FL, lack mechanisms to detect and adapt to changing data distributions, rendering them unfit for addressing fairness under group-specific distributed concept drift. In such scenarios, characteristics or patterns observed in one concept may not be applicable to others, intensifying fairness challenges in FL environments. Hence, using a multi-model approach offers a promising solution to address the complexities of concept drift while preserving fairness across diverse data distributions.
II-C Federated Learning under Concept Drift
Concept drift has been extensively studied in the centralized setting, and readers can refer to the surveys [12, 13] for further details. However, applying centralized algorithms to FL may not be well-suited for handling distributed concept drifts with heterogeneous data distributions across time and clients [14, 15, 16, 17]. In the context of FL, research on concept drift is still in its early stages and most works do not account for the distributed nature of concept drift in their algorithms [14].
In their recent work, Jothimurugesan et al. [14] introduce a novel approach to tackle the challenge of distributed concept drift in FL. They treat drift adaptation as a dynamic clustering problem, emphasizing the limitations of single global models when dealing with distributed drifts. To address this issue, the authors propose a multi-model approach that leverages local drift detection based on the overall loss, offering a more robust solution to adapt to evolving data distributions over time. However, it is important to note that this algorithm, referred to as FedDrift, focuses on the global loss and does not consider group-specific losses. In the context of our work, where we specifically address group-specific distributed concept drift, FedDrift’s reliance on global loss renders it ill-suited for our scenario. Group-specific concept drift may not be effectively detected by FedDrift, as it does not account for group-specific loss variations in local models that are crucial for capturing changes specific to certain groups. This limitation underscores the need for a more nuanced approach to address fairness concerns under group-specific distributed concept drift, which we elaborate on in the subsequent sections.
II-D Limitations and Research Context
While certain studies have delved into fairness under concept drift in centralized streaming environments [18, 19, 20, 21], it is important to recognize their limitations when applied to fairness under concept drift in FL. These fairness-aware algorithms may not be well-suited for FL due to several reasons. Firstly, they typically require direct access to sensitive attributes, which is restricted in FL where only model parameters are shared among clients and the data of each client is kept private. Secondly, these algorithms rely on decision trees which may pose a security risk in a FL setting when sharing decision tree models with the server, requiring enhanced privacy and security measures. In addition, these studies primarily operate in a streaming context, with one example arriving at a time, which differs from our setup that is characterised by the simultaneous processing of multiple examples, and contrasts with the typical use of neural networks in FL.
Finally and most importantly, these existing works primarily concentrate on the study of concept drift affecting both groups simultaneously, rather than delving into datasets characterized by group-specific concept drift or its distributed counterpart. This limitation constrains the assessment of their approaches under realistic conditions with evolving fairness dynamics, where different groups may undergo distinct patterns of concept drift. In this work, we endeavor to bridge this gap by pioneering the formalization and exploration of group-specific concept drift in the context of FL, unveiling its importance in advancing fairness in machine learning.
III Problem Statement
We first define the following notations used throughout the paper. We assume there exist $K$ clients and $T$ timesteps in the FL setting. In our context, a timestep refers to a specific interval during which new data is received by a client, which is composed of $R$ communication rounds. The collection of private data received by a client $k$ at timestep $t$ is represented as $D_{k}^{t}\{X,Y\},k\in K,t\in T$. Each dataset $D_{k}^{t}$ contains $|D_{k}^{t}|$ instances. We further assume a classification setting, where $X$ is the input space, $Y$ is the output space, $\hat{Y}$ represents the predicted class, and $S$ is the sensitive attribute of $X$.
III-A Classical Federated Learning
The primary goal of FL is to train a central model, $\theta$, located on the server, while preserving the privacy of the clients’ data. In FL, multiple clients collaborate to train a model that minimizes the weighted average of the loss across all clients. The objective of this framework can be expressed as follows, as originally proposed in [3]:
$\min_{\theta}f(\theta)=\sum_{k\in K}\frac{|D_{k}|}{|D|}G_{k}(\theta),\qquad G_{k}(\theta)=\frac{1}{|D_{k}|}\sum_{i=1}^{n_{k}}f_{i}(\theta),$ (1)
where $G_{k}(\theta)$ represents the local objective for client $k$, and $f_{i}(\theta)$ represents the loss of data point $i$ from client $k$. It is important to note that this classical formulation does not explicitly account for the temporal variance of data across different timesteps.
This original work [3] led to the proposal of an algorithm called Federated Averaging (FedAvg) which involves periodically averaging the locally trained models of clients at each communication round to generate the global model. The FedAvg global model at timestep $t$ can be mathematically expressed as:
$\theta^{t}=\sum_{k\in K}\frac{|D_{k}|}{|D|}\theta^{t}_{k}$ (2)
This equation represents the aggregation of the locally updated models weighted by the ratio of their respective dataset sizes.
Figure 1: Scenarios of concept drift in Federated Learning: 0 - no concept drift; 1 - general concept drift; 2 - distributed concept drift; 3 - group-specific general concept drift; 4 - group-specific distributed concept drift.
III-B Concept Drift
We assume that the underlying distribution of each $D_{k}$ is not constant over time. This means that the characteristics of the data can vary across time, resulting in concept drifts. Concept drifts refer to changes in the joint distribution, where the distribution $P^{t}(X,y)$ is different from $P^{t-1}(X,y)$ at two different timesteps [12]. We are particularly interested in observing real concept drifts, which occur when the distribution $P^{t}(y|X)$ is different from $P^{t-1}(y|X)$ and require updating the model.
In this work, we introduce and formalize two distinct types of concept drifts: group-specific concept drift, and group-specific distributed concept drift.
Definition 1 (Group-Specific Concept Drift)
Let $P^{t}(y|X,S)$ represent the conditional distribution of $y$ given $X$ and group $S$ at timestep $t$. We first define group-specific concept drift as follows:
  • For group $S=s$, there is a concept drift if $P^{t}(y|X,S=s)\neq P^{t-1}(y|X,S=s)$ for any $t>1$.
  • For group $S\neq s$, there is no concept drift if $P^{t}(y|X,S\neq s)=P^{t-1}(y|X,S\neq s)$ for any $t>1$.
In other words, group-specific concept drift occurs when one group experiences a change in the conditional distribution of the target variable over time, while other groups’ conditional distribution remain constant. This difference in concept drift between groups can result in a decrease in fairness, even if loss remains fairly stable. This occurs due to changes in loss within group $S=s$ attributed to concept drift that may not be readily noticeable in the overall loss, particularly when there is an imbalance in the dataset with respect to the sensitive attribute (i.e. when group $S=s$ is underrepresented).
In these situations, conventional drift detection mechanisms that rely solely on overall loss metrics may struggle to capture these underlying dynamics. This underscores the need to incorporate group-specific loss monitoring over time, as fairness concerns can arise when shifts in concept drift affect opportunities for favorable outcomes across groups. It is important to note that, although studies have delved into fairness under concept drift in centralized environments, the concept of group-specific concept drift had not been formally addressed in the literature. In addition, empirical experiments to validate its impact are lacking as the literature focuses on general concept drift, overlooking the nuances introduced by group-specific variations.
Definition 2 (Group-Specific Distributed Concept Drift)
Let $P^{t}_{k}(y|X,S)$ represent the conditional distribution of $y$ given $X$ for client $k$ at timestep $t$ and group $S$. In FL, we first define group-specific distributed concept drift as follows:
  • For each client $k$, there can be a group-specific concept drift i.e., $P^{t}_{k}(y|X,S=s)\neq P^{t-1}_{k}(y|X,S=s)$ and $P^{t}_{k}(y|X,S\neq s)=P^{t-1}_{k}(y|X,S\neq s)$ for any $t>1$.
  • The conditional distribution for client $k$ may differ from or be the same as the conditional distribution for client $k^{\prime}$ at either the same timestep $t$ or a different timestep $t^{\prime}$ i.e. $P^{t}_{k}(y|X,S)=P^{t^{\prime}}_{k^{\prime}}(y|X,S)\ \text{or}\ P^{t}_{k}(y|X,S)\neq P^{t^{\prime}}_{k^{\prime}}(y|X,S)$, where $(t=t^{\prime}\ \text{or}\ t\neq t^{\prime})$.
In other words, group-specific distributed concept drift occurs when different clients in a FL setting experience either distinct or similar group-specific concept drifts that can happen at the same or different timesteps and clients. These temporal and spacial dynamics can lead to challenges in maintaining fairness and accuracy over time.
It is important to highlight that the exploration and comprehensive understanding of group-specific distributed concept drift in the context of FL has not been introduced and explored in the existing literature. The unique challenges posed by this phenomenon, including the temporal misalignment and spatial variations in group-specific concept drift across different clients and timesteps, add an additional layer of complexity to the already intricate landscape of fairness in FL. As such, our work contributes to filling this gap by formally introducing and addressing the intricacies of group-specific distributed concept drift, shedding light on its potential impact on fairness within FL systems.
Figure 1 illustrates five different concept drift scenarios that can arise in FL with three distinct concepts (‘A’, ‘B’ and ‘C’). We focus on studying scenario 4 - group-specific distributed concept drift, as it presents a unique and unexplored aspect of concept drift in the FL setting. In this scenario it can be observed that group-specific concepts differ across time and across clients. By understanding and addressing the intricacies of this scenario, we aim to enhance the fairness of federated models when clients and sensitive attributes experience group-specific distributed concept drifts.
III-C Objective
We present the optimization objective employed to detect and address group-specific distributed concept drift while upholding the privacy of participating clients’ data within the federation. We adopt a multiple global model approach where we aim to discover clusters of clients, each representing a distinct concept, and train each global model accordingly. To achieve this, we incorporate local group-specific loss monitoring on each client. In other words, for each group $s\in S$ we calculate its loss on each global model $\theta_{m}$ in the set of available global models, $GM$, and assign each client to a specific global model, accordingly.
This idea is inspired by previous work [14], which adopts a multiple global model approach based on global loss monitoring. However, the use of global loss monitoring in their approach poses limitations when dealing with group-specific concept drift. This is because global loss may not vary significantly over time in the presence of group-specific concept drift, especially when the dataset is imbalanced with respect to a sensitive group. In this case, changes in loss predominantly occur within specific groups. Consequently, relying solely on global loss may overlook variations introduced by group-specific concept drift.
We employ the following optimization objective for each global model $\theta_{m}$ in the set of global models, $GM$, at each timestep $t$:
$\begin{split}\min_{\theta_{m}\in GM}f(\theta_{m})^{t}=\frac{\sum_{k\in K}G_{k}(\theta)\sum_{t^{\prime}\leq t:w_{k}^{t^{\prime}}=m}|D_{k}^{t^{\prime}}|}{\sum_{k\in K}\sum_{t^{\prime}\leq t:w_{k}^{t^{\prime}}=m}|D_{k}^{t^{\prime}}|}\\ w_{k}^{t}=\arg\min_{m}\sum_{s\in S}\ell(\theta_{m})^{D_{k}^{t}}_{s}\\ \text{s.t.}\quad\forall\theta_{m}\in GM:\forall k:w_{k}^{t}=m:\forall s\in S:\\ \ell(\theta_{m})^{D_{k}^{t}}_{s}\leq\ell(\theta_{m})^{D_{k}^{t-1}}_{s}+\delta_{s}\end{split}$ (3)
where:
  • $t\in T$ denotes the timestep;
  • $k\in K$ represents a client within the federation;
  • $s\in S$ refers to a sensitive group;
  • $\theta_{m}\in GM$ refers to a specific global model;
  • $t^{\prime}\leq t:w_{k}^{t^{\prime}}=m$ is the set of timestamps where client $k$ has identified with global model $\theta_{m}$;
  • $f(\theta_{m})^{t}$ is the weighted sum of local objectives $G_{k}(\theta)$ of clients that identify with global model $\theta_{m}$;
  • $\ell(\theta_{m})^{D_{k}^{t}}_{s}$ is the loss for group $s$ and client $k$ on model $\theta_{m}$ at timestep $t$;
  • $\delta_{s}$ represents a predefined threshold for the acceptable loss difference of group $s$ between two consecutive timestamps on a global model  [^1] As discussed in [14], every client $k$ preserves its entire history of client identities $w_{k}^{t^{\prime}\leq t}$ and local datasets at each timestamp $D_{k}^{t^{\prime}\leq t}$. To mitigate this overhead, clients could opt to retain only a sliding window of the latest timesteps instead.;
Hence, at each timestep, each client identities with the model that yields the lowest sum of group losses, as long as the loss difference of each group $s$ between the current and the previous timestamp does not surpass the threshold for that group, $\delta_{s}$. If there is no such model, then a new global model will be created and added to the set of global models $GM$. Further details will be explained in the next section.
IV FairFedDrift: Fair Federated Learning under Group-Specific Distributed Concept Drift
Our work offers a solution to address the challenge of group-specific distributed concept drift building upon an existing algorithm, FedDrift, presented in [14]. Algorithm 1 presents the pseudo-code of FairFedDrift.  [^2] Source code can be found at: https://github.com/teresalazar13/FairFedDrift.
Algorithm 1 FairFedDrift
1:  Initialize global models GM: $\left[\theta_{0}\right]$, number of timesteps: $T$, number of rounds: $R$, number of clients: $K$, client identities $w_{k\in K}^{t=0}=0$, local epochs: $E$, local batch size: $|B|$, local learning rate: $\eta$.
2:  for each timestep $t\in T$ do
3:     for each client $k\in K$ do
4:        if $\nexists\theta_{m}\in GM:\forall s\in S,\ell(\theta_{m})^{D_{k}^{t}}_{s}\leq\ell(\theta_{m})^{D_{k}^{t-1}}_{s}+\delta$ then
5:           Add new $\theta$ to GM at index $i$
6:           $w_{k}^{t}=i$
7:        else
8:           $w_{k}^{t}=\arg\min_{m}$
       $\sum_{s\in S}\begin{cases}\ell(\theta_{m})^{D_{k}^{t}}_{s}\\ \quad\quad\text{if }\ell(\theta_{m})^{D_{k}^{t}}_{s}\leq\ell(\theta_{m})^{D_{k}^{t-1}}_{s}+\delta_{s}\\ \infty\quad\text{else}\end{cases}$
9:        end if
10:     end for
11:     for each $i,j$ in parallel from 1 to $|GM|$ do
12:        $L_{ij}-L_{ii}\leftarrow\max_{k:\exists t^{\prime}<t:w_{k}^{t^{\prime}}=j,l:\exists t*<t:w_{l}^{t*}=i}$
       $\sum_{s\in S}\begin{cases}\ell(\theta_{i})^{\cup_{t^{\prime}}D_{k}^{t^{\prime}}}_{s}-\ell(\theta_{i})^{\cup_{t*}D_{l}^{t*}}_{s}\\ \quad\quad\text{if }\ell(\theta_{i})^{\cup_{t^{\prime}}D_{k}^{t^{\prime}}}_{s}-\ell(\theta_{i})^{\cup_{t*}D_{l}^{t*}}_{s}\leq\delta_{s}\\ \infty\quad\text{else}\\ \end{cases}$
13:        Compute $L_{ji}-L_{jj}$ similarly
14:        $Z_{ij}$ = $\max(L_{ij}-L_{ii},L_{ji}-L_{jj},0)$
15:     end for
16:     while $\min Z_{ij}\neq\infty$ do
17:        Add new model to $GM$: $\theta_{k}=\frac{\theta_{i}W_{i}+\theta_{j}W_{j}}{W_{i}+W_{j}}$,
       $W_{n\in i,j}=\sum_{k\in K}\sum_{t^{\prime}<t:w_{k}^{t^{\prime}}=n}|D_{k}^{t^{\prime}}|$
18:        $Z_{kl}=\max(Z_{il},Z_{jl},0)$ for all $l$
19:        $w_{k}^{t}$ = $w_{i}^{t}$ + $w_{j}^{t}$
20:        Delete $\theta_{i}$ and $\theta_{j}$
21:     end while
22:     for each round $r\in R$ do
23:        for each client $k\in K$ do
24:           for each global model $\theta_{m}\in GM$ do
25:              $\theta_{m,k}$ = $\theta_{m}$
26:              for each local epoch $e\in E$ do
27:                 for batch $b\in B$ from $\cup_{t^{\prime}\leq t:w_{k}^{t^{\prime}}=m}D_{k}^{t^{\prime}}$ do
28:                    $\theta_{m,k}$ = $\theta_{m,k}-\eta\nabla\ell(\theta_{m},b)$
29:                 end for
30:              end for
31:           end for
32:        end for
33:        for each global model $\theta_{m}\in GM$ do
34:           $\theta_{m}$ = $\frac{\sum_{k\in K}\theta_{m,k}\sum_{t^{\prime}\leq t:w_{k}^{t^{\prime}}=m}|D_{k}^{t^{\prime}}|}{\sum_{k\in K}\sum_{t^{\prime}\leq t:w_{k}^{t^{\prime}}=m}|D_{k}^{t^{\prime}}|}$
35:        end for
36:     end for
37:  end for
IV-1 Multi-model approach
The core idea behind this algorithm is to recognize that a single global model is ill-suited for handling distributed group-specific concept drifts, leading to the adoption of a multi-model strategy. In the beginning of the federation, there is a single global model, $\theta_{0}$, and all clients are associated with it, $w_{k\in K}^{t=0}=0$. At each timestep, $t$, each client $k$ has new incoming data $D_{k}^{t}$ and associates itself with the model that yields the lowest sum of losses of every group $s\in S$, as long as the difference in loss of each group $s$ between the current and the previous timestep does not surpass the threshold for that group, $\delta_{s}$. In contrast to FairFedDrift, which is designed with fairness considerations, FedDrift is fairness-unaware and associates itself with the model that yields the lowest global loss, without accounting for group-specific losses.
IV-2 Group-specific distributed concept drift detection:
A drift is identified if there is no global model in the set of global models where for any group $s\in S$ the difference in losses of two consecutive timestamps of that group is smaller than the threshold for that group, $\delta_{s}$ (line 4). In this case, a new global model is created. Contrary to FairFedDrift, in FedDrift a drift is identified if the difference between the lowest global loss (not group-specific) of two consecutive timestamps is greater than a threshold, $\delta$.
IV-3 Global model merging
Importantly, since new concepts might simultaneously emerge at different clients, a merge operation is employed to consolidate models that correspond to the same concept. This operation involves the creation of a distance matrix (lines 11-15), $Z$, where each cell is calculated for each pair of global models, $i$ and $j$. The distance $Z_{ij}$ corresponds to the maximum value of $L_{ij}-L_{ii}$, $L_{ji}-L_{jj}$, and 0, where $L_{ij}-L_{ii}$ represents the sum of the group loss differences of model $\theta_{i}$ when tested on data associated with model $j$ and $i$. For a model $j$, there can be multiple clients $k$ that have identified with it: $k:\exists t^{\prime}<t:w_{k}^{t^{\prime}}=j$. Here, $\cup_{t^{\prime}}D_{k}^{t^{\prime}}$ represents the data of client $k$ associated with model $j$. If the loss difference of any group $s$ exceeds the threshold for that group, $\delta_{s}$, the value $\infty$ is assigned to $Z_{ij}$. The merging process (lines 16-21) combines the two models with the lowest distances in the matrix (provided they do not reach $\infty$) and unifies cluster identities by averaging their models, with the weighting based on the size of each model’s training dataset. The distance between clusters of multiple elements is measured as the maximum distance between their constituent parts. In contrast to FairFedDrift, FedDrift only focuses on the global loss for merging global models, without considering group-specific losses.
IV-4 Training and Model Averaging
The training and model averaging process (lines 22-36) involves multiple communication rounds, denoted by $R$. In each round, every global model $\theta_{m}$ in $GM$ is updated based on the contributions from clients associated with that model. The update is performed through iterative optimization, where each client $k$ trains its local model $\theta_{m,k}$ if it had been associated with model $m$ until that timestep. These local models are then aggregated to update the global model $\theta_{m}$ by considering the weighted sum of individual models, where the weights are proportional to the sizes of the respective client datasets.
To sum up, the choice of considering the loss of each group $s\in S$ in FairFedDrift, in contrast to FedDrift [14], stems from its effectiveness in handling group-specific concept drift as it inherently better suited to capture the nuances of fairness in these scenarios where specific groups may experience concept drift at different timesteps. This is attributed to the fact that, in instances of group-specific concept drift, a global loss metric may not manifest substantial variations. We elaborate on this aspect in our experiments section, providing a more comprehensive understanding of the dynamics involved.
V Methodology and Experimental Design
In our experimental setup, we operate within a FL framework that operates continuously over time involving K clients. At each timestep, every client receives a new batch of training data. After completing training for a given timestep, which can span multiple communication rounds, for all clients, we test the global model associated with that client over the batch of data arriving at that client at the following timestep (between lines 3 and 4, in Algorithm 1).
V-A Datasets and Distributed Group-Specific Concept Drift Simulation
In our experimental setup, we leverage diverse datasets derived from established benchmarks in machine learning: MNIST [22], FEMNIST [23], and Adult [24]. These datasets were chosen due to their extensive size, making them well-suited for this FL setup, where each client necessitates a substantial number of data instances across different timesteps. In addition, these datasets have also been employed in fairness studies [25, 26], which makes them suitable to our experiments.
To simulate this FL scenario, we partition each dataset into $K$ clients and $T$ timesteps. Furthermore, we introduce group-specific concept drift within the datasets at specific timesteps, enabling the evaluation of the efficacy of our proposed method under evolving data conditions.
V-A1 MNIST-GDrift
The MNIST dataset consists of gray-scale handwritten digit images, commonly used for digit recognition tasks [22]. We introduce one sensitive attribute, $S$, with two groups ($S=1$ and $S=0$), representing distinct image characteristics. For $S=0$ images, we invert the background and digit colors in comparison to standard MNIST images ($S=1$). To simulate two distinctive group-specific concepts (‘B’ and ‘C’) from the original distribution (‘A’), we label-swap images for $S=0$: for drift ’B’, we swap labels ‘0’ and ‘1’; and for drift ‘C’, we swap labels ‘1’ and ‘2’.
V-A2 FEMNIST-GDrift
The FEMNIST is a dataset featuring gray-scale images of 10 types of clothing items, designed as an alternative to MNIST for benchmarking image classification algorithms [23]. To create different groups and simulate group-specific concept drift, we use the same strategies as the MNIST-GDrift dataset.
V-A3 Adult-GDrift
We experiment on the Adult dataset [24] commonly used in fair machine learning research. The prediction task for the Adult dataset involves determining whether a person’s income exceeds \$50K/yr. In this context, race serves as the sensitive attribute, where $S=1$ typically corresponds to White individuals, and $S=0$ represents non-White individuals. To introduce concept drift, we modify features of the samples where $S=0$. For concept ‘A’, the dataset remains unchanged. For concept ‘B’, the target class of non-White individuals who are Male is set to 1 (positive). For concept ‘C’, the target class of non-White individuals whose ‘relationship’ feature is not ‘Husband’ is set to 1.
V-A4 $\alpha$ parameter
The parameter $\alpha$ is a crucial element in our experimental design, determining the size of the unprivileged group relative to the privileged group. The size of the unprivileged group is a result of multiplying the size of the privileged group by $\alpha$. By adjusting the value of $\alpha$, we can control the balance between the groups in our datasets, allowing us to explore different scenarios where one group may be underrepresented or over-represented relative to the other. As such, when $\alpha$ is closer to zero, the relative size of the unprivileged group ($S=0$) is smaller. To generate a specific number of instances depending on the value of $\alpha$ for the MNIST-GDrift and FEMNIST-GDrift datasets, we adjust the number of transformed images accordingly. However, for the Adult-GDrift dataset, the process is more intricate as the original dataset is oversampled, whose default representation is equivalent to $\alpha=0.16$. If $\alpha>0.16$, we generate privileged instances, and if $\alpha<0.16$, we generate unprivileged instances. We employ SMOTE [27] for creating these instances.
V-B Implementation Details
We employ a convolutional neural network (CNN) on the MNIST-GDrift and FEMNIST-GDrift datasets featuring a convolutional layer, a max-pooling layer, a dense layer, and a softmax activation function (multi-class classification). For the Adult-GDrift dataset we employ a fully connected feed-forward neural network with one hidden layer with 10 $tanh$ units, and a sigmoid output neuron (binary classification). We set $K=10$, $T=10$, $\eta=0.1$, and experiment with different values of $\delta_{s}$: $\{0.05,0.1,0.25,0.5,0.75,1.0,1.25,1.5\}$. Furthermore, for the MNIST-GDrift and FEMNIST-GDrift datasets we set $R=10$, $|B|=32$, $E=5$. For the Adult-GDrift dataset we set $R=1$, $|B|=10$, $E=10$.
V-C Metrics
We evaluate the algorithms based on both accuracy (ACC) and fairness (AEQ, OEQ, OPP). Accuracy Equality (AEQ) [28] is a measure of fairness that calculates the ratio of accuracies across two groups ($S=0$ and $S=1$):
Definition 3 (Accuracy Equality (AEQ))
Accuracy Equality (AEQ) [28] is a measure of fairness that calculate the ratio of accuracies across different groups:
$\text{AEQ}=\frac{P[\hat{Y}=Y\mid S=0]}{P[\hat{Y}=Y\mid S=1]}$ (4)
Definition 4 (Overall Equality of Opportunity (OEQ))
Overall Equality of Opportunity (OEQ) [29] calculates the ratio of the average true positive rates (TPRs) across all target classes for two groups. Let $P[\hat{Y}=k\mid S=0,Y=k]$ and $P[\hat{Y}=k\mid S=1,Y=k]$ represent the TPRs for group $S=0$ and $S=1$ for the $k^{th}$ target class, respectively. The formula for OEQ is given by:
$\text{OEQ}=\frac{1}{K}\sum_{k=1}^{K}\frac{P[\hat{Y}=k\mid S=0,Y=k]}{P[\hat{Y}=k\mid S=1,Y=k]}$ (5)
where $K$ is the number of target classes.
Definition 5 (Overall Predictive Parity (OPP))
Overall Predictive Parity (OPP) [30] calculates the ratio of the average positive predictive values (PPVs) across all target classes for two groups. Let $P[Y=k\mid S=0,\hat{Y}=k]$ and $P[Y=k\mid S=1,\hat{Y}=k]$ represent the PPVs for group $S=0$ and $S=1$ for the $k^{th}$ target class, respectively. The formula for OPP is given by:
$\text{OPP}=\frac{1}{K}\sum_{k=1}^{K}\frac{P[Y=k\mid S=0,\hat{Y}=k]}{P[Y=k\mid S=1,\hat{Y}=k]}$ (6)
where $K$ is the number of target classes.
In these metrics (AEQ, OEQ, OPP), a value of 1 indicates ideal fairness, and the numerator corresponds to the group with the lowest value to account for cases of reverse discrimination.
Figure 2: Test Scenarios of Group-Specific Distributed Concept Drift.
V-D Group-Specific Distributed Concept Drift Scenarios
For our experiments, we design three distinct scenarios of group-specific distributed concept drift, as illustrated in Figure 2. These aim to provide a comprehensive exploration of the temporal and spatial dynamics that can impact fairness and accuracy in FL settings.
V-D1 Baselines
FairFedDrift is compared against three pivotal algorithms: FedAvg, FedDrift [14] and Oracle. The Oracle algorithm serves as an idealized benchmark, equipped with access to concept IDs during training and employing multiple-model training based on ground-truth clustering. Given the unique problem of group-specific distributed concept drift, traditional baselines for fair FL are absent, as explained in the Related Work section. We found that these achieve very similar results to FedAvg on these datasets under group-specific distributed concept drift, as they lack the adaptability to effectively address the challenges posed by these evolving data distributions. Nevertheless, Oracle serves as a perfect algorithm for this problem setup and should act as an ideal baseline.
VI Experimental Results and Discussion
In this section, we unveil the outcomes of our experiments, addressing a distinctive challenge in our investigation - that of group-specific distributed concept drift. It is essential to underscore that our contribution extends beyond the mere presentation of algorithms’ results; we introduce and explore this novel problem, shedding light on the intricate challenges it presents.
VI-A Results Across Multiple Datasets and Scenarios
We delve into the evaluation of the algorithms on the described datasets, considering different $\alpha$ values and distinct scenarios of group-specific distributed concept drift presented before. Table I provides the average results of all clients on all timesteps. The results demonstrate the efficacy of FairFedDrift in achieving favorable fairness outcomes (AEQ, OEQ, OPP) with minimal impact on accuracy. Notably, FairFedDrift’s fairness results results closely approaches that of the idealized Oracle algorithm.
The following sections will elucidate the mechanisms behind FairFedDrift’s superior performance and provide an in-depth analysis of why it achieves better fairness outcomes.
Dataset $\alpha$ Metric Algorithm
FedAvg FedDrift FairFedDrift Oracle
MNIST
GDrift
0.1 AEQ 0.84+-0.04 0.86+-0.06 0.92+-0.04 0.93+-0.02
OEQ 0.81+-0.04 0.83+-0.05 0.88+-0.04 0.90+-0.03
OPP 0.80+-0.04 0.83+-0.06 0.88+-0.04 0.90+-0.02
ACC 0.95+-0.01 0.95+-0.02 0.94+-0.02 0.95+-0.01
0.05 AEQ 0.81+-0.04 0.84+-0.05 0.87+-0.06 0.90+-0.04
OEQ 0.77+-0.03 0.80+-0.05 0.81+-0.06 0.84+-0.05
OPP 0.76+-0.04 0.79+-0.05 0.81+-0.07 0.84+-0.04
ACC 0.95+-0.01 0.96+-0.01 0.95+-0.02 0.95+-0.01
FEMNIST
GDrift
0.1 AEQ 0.83+-0.05 0.84+-0.05 0.90+-0.02 0.91+-0.01
OEQ 0.77+-0.03 0.77+-0.05 0.82+-0.04 0.84+-0.02
OPP 0.77+-0.04 0.77+-0.05 0.82+-0.03 0.84+-0.02
ACC 0.87+-0.01 0.85+-0.02 0.86+-0.02 0.87+-0.02
0.05 AEQ 0.78+-0.07 0.81+-0.07 0.84+-0.05 0.86+-0.04
OEQ 0.68+-0.05 0.70+-0.05 0.72+-0.06 0.74+-0.05
OPP 0.68+-0.06 0.69+-0.07 0.72+-0.06 0.74+-0.05
ACC 0.88+-0.02 0.89+-0.02 0.87+-0.02 0.87+-0.02
Adult
GDrift
0.1 AEQ 0.83+-0.11 0.90+-0.04 0.91+-0.02 0.91+-0.02
OEQ 0.63+-0.19 0.72+-0.15 0.67+-0.15 0.68+-0.15
OPP 0.70+-0.08 0.73+-0.11 0.77+-0.10 0.77+-0.10
ACC 0.80+-0.01 0.81+-0.01 0.81+-0.02 0.82+-0.02
0.05 AEQ 0.82+-0.13 0.87+-0.08 0.91+-0.02 0.91+-0.02
OEQ 0.63+-0.21 0.72+-0.14 0.69+-0.15 0.70+-0.14
OPP 0.69+-0.10 0.73+-0.10 0.77+-0.11 0.78+-0.10
ACC 0.81+-0.01 0.82+-0.01 0.82+-0.02 0.82+-0.01
TABLE I: Average results for the MNIST-GDrift, FEMNIST-GDrift, and Adult-GDrift datasets under different $\alpha$ values.
VI-B Unveiling the Influence of $\alpha$ on Loss Over Time
Figure 3: Average clients’ loss ($\ell$) and group-specific ($S=0$) loss ($\ell_{S=0}$) results for the MNIST-GDrift dataset considering $\alpha=0.1$ and $\alpha=0.5$.
Figure 4: Clustering of FedDrift, FairFedDrift and Oracle on the MNIST-GDrift dataset, considering $\alpha=0.1$. Each cell represents the model ID at each client and timestep, while background colors denote the ground-truth concept.
We delve into the influence of group-specific distributed concept drift on the loss, emphasizing scenarios where the overall loss remains relatively stable while the loss of the unprivileged group experiences a noticeable decrease.
Figure 3 illustrates the loss ($\ell$) and group-specific ($S=0$) loss ($\ell_{S=0}$) results obtained from running the FedAvg algorithm on the MNIST-GDrift dataset considering $\alpha=0.1$ and $\alpha=0.5$. We showcase the results for the FedAvg since both FedDrift and FairFedDrift are built on top of this base algorithm.
The results for $\alpha=0.1$ reveal that the loss of the unprivileged group varies significantly over time while the overall loss remains relatively stable. In such scenarios, FedDrift can struggle to detect significant loss changes, making the tuning of $\delta$ challenging since smaller values can result in too many clusters, while larger values may fail to detect group-specific concept drift. This highlights the limitations of FedDrift in addressing subtle, group-specific changes, as demonstrated before.
Conversely, when $\alpha=0.5$, both the overall loss and the loss of the unprivileged group exhibit noticeable variations. In these cases, FedDrift can successfully detect loss changes and create clusters accordingly, showcasing its capability in scenarios where the unprivileged group is well-represented and its losses are integral indicators of concept drift.
VI-C Comparative Cluster Creation: FedDrift vs. FairFedDrift
Figure 4 presents the clustering outcomes achieved by FedDrift, FairFedDrift and Oracle on the MNIST-GDrift dataset considering $\alpha=0.1$. Although both FedDrift and FairFedDrift can detect concept drift in the third timestep, it can be observed in the fourth timestep that clients in the green concept in FedDrift prefer to associate themselves with the original model (‘0’), while FairFedDrift’s clients in the green concept associate themselves with models created in the previous timestep. While FairFedDrift’s approach to associating clients according to the sum of losses (line 8 in Algorithm 1) is crucial in this scenario, FedDrift lacks this mechanism, resulting in clients preferring the original model. In addition, it can be observed that FairFedDrift successfully merges models corresponding to the same concept (e.g. ‘1-3-4-5’). The visual analysis underscores FairFedDrift’s ability to achieve more accurate and adaptive clustering, thereby offering a robust solution this problem.
Finally, FairFedDrift offers easier tuning in situations of group-specific concept drift compared to FedDrift. The loss of the unprivileged group undergoes significant variations over time, while the overall loss remains relatively stable. This makes tuning FedDrift more challenging, as it may struggle to detect significant loss changes, making the tuning of its overall loss threshold problematic – smaller values result in excessive clusters, while larger values may fail to detect group-specific concept drift altogether.
VI-D Effect of Window Size on Fairness and Performance
Finally, we explore the influence of window size on both fairness and performance metrics in our FL framework. We examine how varying window sizes affect the system’s ability to adapt to concept drift while considering the trade-offs between overhead and model effectiveness. To evaluate the impact of window size, we conducted experiments across different window size configurations, ranging from retaining the entire history of client identities and local datasets to utilizing only a sliding window of the latest timesteps.
Figure 5 illustrates the results of this experiment. It can be observed that varying the window size had minimal effect on both fairness and performance across different configurations. Despite the reduction in historical data retention, the system maintained comparable levels of fairness and performance across varying window sizes.
Given the negligible impact of window size on fairness and performance, it is possible to mitigate the overhead associated with preserving the entire history of client identities and local datasets - clients can opt to maintain only a sliding window of the latest timesteps. By adopting this approach, clients can significantly reduce the computational and storage burden without compromising fairness or performance. Moreover, in certain scenarios, forgetting older data might actively aid in adaptating to concept drift, enhancing the system’s resilience to changing data distributions over time.
Figure 5: Effect of window size on fairness and performance on the MNIST-GDrift dataset considering $\alpha=0.1$.
VII Conclusions and Future Work
This work presents a pioneering effort in introducing and formalizing the problem of group-specific distributed concept drift within FL frameworks. Our experiments underscore the detrimental impact of this phenomenon on fairness, emphasizing the need for effective mechanisms to address evolving data distributions while upholding fairness principles. Our contributions include the formalization of group-specific concept drift and its distributed counterpart, an experimental framework for studying it, and the application of model clustering and fairness monitoring techniques to mitigate the challenges posed by this issue.
We demonstrate that in the absence of group-specific distributed concept drift handling, fairness tends to decrease over time. In addition, for scenarios with high group imbalance, our findings indicate that overall loss and accuracy do not change significantly. However, there is a noticeable decrease in fairness due to the changing loss of the unprivileged group. In contrast, as the representation of the unprivileged group increases, both overall loss and the loss of the unprivileged group exhibit more noticeable changes. Our proposed algorithm, FairFedDrift, emerged as a practical solution for achieving fairness in FL amid group-specific distributed concept drift. By continuously monitoring group loss and adapting to changes in data distributions, FairFedDrift demonstrated its effectiveness in mitigating unfairness.
As we look to the future, several directions for further research emerge. Automating the hyperparameter selection process for FairFedDrift could streamline its application across diverse scenarios. Additionally, developing specialized fairness metrics tailored for evaluating concept drift in distributed learning settings would provide more nuanced insights into the evolving landscape of fairness.
The intricate issues of fairness and group-specific distributed concept drift demand ongoing exploration and refinement. Our work lays the foundation for addressing these challenges as machine learning continues to evolve in sensitive domains.
Ethical Implications
While our research aims to achieve fairness, it is essential to recognize that this objective may not be guaranteed in all use cases. The fairness metrics selected for evaluation and the choice of sensitive attributes are often contingent upon the application domain. Moreover, the binarization process employed to represent different groups of sensitive attributes may not capture the diversity within each group. These considerations underscore the complexity of fairness in machine learning and highlight the importance of ongoing refinement to address evolving ethical concerns.